# How does the entanglement entropy of a many-body quantum system change after a single measurement?

### v3

@@ Abstract
For one-dimensional non-interacting complex fermions, we compute numerically the probability distribution of the change in the entanglement entropy (EE), after saturation, resulting from a single measurement of the occupation number by using different measurements protocols. In the thermodynamic limit, the system is in the area-law phase for any monitoring strength, however we can also study the distribution in the critical phase characterized by a logarithmic scaling of the EE with system size by considering a sufficiently weak monitoring strength. For the quantum jump and the projective measurement protocols, we observe clear deviations from Gaussianity characterized by broader and asymmetric tails, exponential for positive values of the change, and a peak at zero, likely a precursor of Zeno effect, that increases with the system size and the monitoring strength. Intriguingly, the distribution is spatially inhomogeneous. For sites around the boundary separating the subsystems defining the EE, the distribution is close to Gaussian with a broad support while for the rest of sites has asymmetric exponential tails and a much narrower support. As the monitoring strength increases, the full distribution is controlled by the boundary sites. For a quantum state diffusion protocol, where the change in EE is defined between two consecutive time steps, the distribution, Gaussian for weak monitoring, gradually develops symmetric exponential tails. For strong monitoring, the core turns from Gaussian to strongly peaked at zero suggesting the dominance of quantum Zeno effect.
@@ I Introduction
The quantum measurement problem has gained new impetus in recent times not only because the availability of the mentioned experimental measurement set-ups with a high level of control but also because for quantum computation, and other quantum technologies, it is fundamental to extract information from quantum states with a minimal level of disturbance on the system so that quantum coherence is preserved. This is a challenging task as repeated measurements can trigger [23, 24, 25, 26, 27, 21, 28, 29, 30, 31, 32, 33, 34] a measurement induced phase transitions (MIPT) in the entanglement entropy, and other quantum information observables non-linear in the density matrix, separating phases with different scaling with respect to the system size. Therefore, a high monitoring rate has the potential of destroying quantum features that may be necessary for practical applications. Although challenging because of the post-selection problem, the existence of MIPT has already [35, 36] been confirmed experimentally.
@@
The precise characterization of the MIPT in interacting many-body quantum systems are still being debated. However, for one dimensional free Dirac fermions, namely, non-relativistic complex fermions, with no disorder and subjected to either continuous [19] or projective [37] measurements there is now growing consensus, despite some initial contradicting results [38, 20], that the system stays in the area-law phase for any monitoring strength. For projective measurements [37], see also [39, 40], this conclusion was reached analytically by using the replica trick and mapping the problem onto a non-linear sigma model [41, 42, 43, 44].
For Majorana fermions, and a continuous monitoring protocol based on weak measurements, the use of a variant of the mentioned non-linear sigma model predicts [45, 40] the existence of a transition at finite monitoring rate. The different symmetries of the associated non-linear sigma model explains the differences between Majorana and Dirac fermions.
Increasing the spatial dimensionality [46], the range of hopping [47, 48, 49] or including interactions [50, 51, 49, 52, 53, 54, 55] seems to induce a MIPT at finite monitoring strength. The interplay between Anderson localization induced by disorder and measurements promises a rich phase diagram which only recently has started to be investigated [56, 57].
In the context of the Sachdev-Ye-Kitaev (SYK) model [58, 59, 60, 61, 62] a setting consisting in two one dimensional coupled chains with an integrable SYK on each site, where the parity is monitored, was shown [63] to experience a MIPT transition analytically for sufficiently large monitoring rate by using the Keldysh path-integral formalism. In the gravity dual, monitoring, at least phenomenologically and in the post-select approximation, seems to be related with additional complex sources in the boundary [64, 65]. By contrast, a two-site SYK initially uncoupled but being continuously monitored undergoes [66] a transition to a state whose gravity dual is a traversable wormhole [67] provided that the monitoring rate is sufficiently strong.
@@
All the above findings are based on the study of averaged (over quantum trajectories) quantities. However, it is unclear whether the average is a good indicator of the observable because of the importance of rare events or other features leading to a non-Gaussian distribution with sufficiently broad tails. However, the study of the full distribution of probability of observables like the entanglement entropy [68] or the local particle density and current [69] has just started. Even in the simplest case of a non-interacting fermionic system in one dimension, studied in Refs. [68, 69], it is not yet known the precise form of these distributions for the different measurement protocols, its system size dependence or its role in the characterization of the transitions. More generally, a detailed understanding of the distribution function may help identify a subset, hopefully small, of quantum trajectories that control the average which could alleviate the so-called post-selection problem [70, 71, 72, 73] that makes challenging the experimental observation of MIPT.
@@ II Model and Method
$H=J\sum_{i=1}^{L}\left[c_{i}^{\dagger}c_{i+1}+c_{i+1}^{\dagger}c_{i}\right],$ (1)
where $c_{i}$ and $c_{i}^{\dagger}$, $i=1,2,\cdots L$ are the annihilation and creation operators respectively for fermions at site $i$. $J$ is the hopping strength which for convenience is set to $J=1$. We impose periodic boundary conditions $c_{i}=c_{i+L},c_{i}^{\dagger}=c_{i+L}^{\dagger}$ with $L$ the lattice size which we set to be even. The Hamiltonian commutes with the particle number operator $\hat{N}=\sum_{i=1}^{L}\hat{n}_{i}$ where $\hat{n}_{i}\equiv c^{\dagger}_{i}c_{i}$ is the occupation number operator at site $i$ which has eigenvalues $0,1$. The eigenvalues of $\hat{N}$ are the total number of fermions $N=1,2,\ldots$. We consider the system at half-filling $N=L/2$. The initial state, termed Néel state, in an occupation number basis is $|\psi(t=0)\rangle=|101010\cdots\rangle$.
@@
(2). Quantum jump (QJ) protocol [10, 11, 12, 5, 13], a continuous monitoring protocol usually referred to as photo-detection in quantum optics where the measurement apparatus is always active and the detection, usually termed a quantum jump, occurs with a certain probability that depends on the monitoring strength $\gamma$. In our case, what is detected is an occupied site, $n_{i}=1$. We model the time evolution of the state by the quantum trajectory method [7, 8, 6, 9] characterized by the jump operator $L_{i}=\sqrt{\gamma}\hat{n}_{i}$. Further details are in Appendix A.2.
@@
$|\psi(t)\rangle=\prod_{k=1}^{N}\left[\sum_{j=1}^{L}U_{jk}(t)c_{j}^{\dagger}\right]|\text{vac}\rangle.$ (2)
@@
$D_{ij}\equiv\langle\psi(t)|c_{i}^{\dagger}c_{j}|\psi(t)\rangle$ (4)
@@
$|\psi(t)\rangle=\sum_{\{j\}}U_{j_{1},1}(t)U_{j_{2},2}(t)\cdots U_{j_{N},N}(t)c^{\dagger}_{j_{1}}c^{\dagger}_{j_{2}}\cdots c^{\dagger}_{j_{N}}|\text{vac}\rangle=\sum_{\{j\}}\left[\prod_{i=1}^{N}U_{j_{i},i}(t)c^{\dagger}_{j_{i}}\right]|\text{vac}\rangle$ (5)
then we substitute Eq. 5 into Eq. 4,
$D_{mn}=\sum_{\{i\},\{j\}}\biggl(\prod_{k=1}^{N}U_{i_{k},k}^{*}\biggr)\,\biggl(\prod_{l=1}^{N}U_{j_{l},l}\biggr)\,\langle\text{vac}|c_{i_{N}}\cdots c_{i_{1}}c_{m}^{\dagger}c_{n}c_{j_{1}}^{\dagger}\cdots c_{j_{N}}^{\dagger}|\text{vac}\rangle=\sum_{k=1}^{N}U^{*}_{mk}U_{nk}$ (6)
@@
Our primary focus is the entanglement entropy, $S=\operatorname{Tr}(\rho_{A}\ln\rho_{A})$, where $\rho_{A}$ is the reduced density matrix for the subsystem $A=\{1,2,\dots,\ell\}$ with $\ell\in\{1,2,\dots,L\}$. The correlation matrix $D^{A}$ in sub-system $A$ is given by the $\ell\times\ell$ upper-left block of $D$, with elements defined as $D^{A}_{ij}=\operatorname{Tr}(\rho_{A}\,c_{i}^{\dagger}c_{j})$. Since $\rho_{A}$ is Gaussian, Wick’s theorem allows us to express it in terms of the eigenvalues of $D^{A}$ [74, 75, 76, 77],
$S_{\ell}(t)=-\sum_{i=1}^{\ell}\left(\lambda_{i}(t)\ln(\lambda_{i}(t))+(1-\lambda_{i}(t))\ln(1-\lambda_{i}(t))\right)$ (7)
@@
Strictly speaking, rather than with the distribution of the full EE at those long times, our main interest is the distribution of the EE difference between the state right before and right after a measurement – for homodyne detection this corresponds to one time step. However, for the sake of completeness, we depict in Fig. 2 the probability distribution of the full EE in this long time region. As is observed, in the weak monitoring limit, the distribution is close to Gaussian for the three protocols. Deviations from Gaussianity, characterized by broad tails and a bimodal distribution are qualitatively similar in all protocols, require a strong monitoring rate which we believe corresponds to a region where the quantum Zeno effect [78, 79, 80, 81, 82] is dominant. We comment that similar results for the full EE distribution were obtained earlier [31] for a continuously monitored Ising chain. These results suggest that the distribution function of the full EE seems to be capable of capturing a transition in the EE (if present). As we mentioned in the introduction, this model is in the area-law phase for $L\to\infty$, and any $\gamma>0$, but we still believe that in the range of $L$ that we explore, and for small $\gamma$, our results describe the distribution of a system in a phase where the EE scales with the logarithmic of the system size. In any case, it is promising that within the range of sizes we can explore, the full distribution captures sharp differences between the weak and strong monitoring rates.
@@ III The quantum state diffusion measurement protocol
We initiate our analysis with continuous measurements characterized by the QSD protocol, see Appendix A.1 for details. This protocol, related to homodyne measurements, is characterized by weak measurements of the particle number on each site of the lattice at each time step $dt$. Therefore, by change in entanglement after one measurement, we really mean the change in EE after one time step because, by definition, at a given time, measurements are carried out on all sites at once so the change after a single measurement is not possible. It is natural to study the speed of the entanglement change at each step $dt$ of the dynamical evolution defined as,
@@
where $t$ is sufficiently large so that the EE has already reached its saturation value and $S_{A}$ is computed using Eq. (7). Results for the probability distribution of $\delta S_{qsd}$, depicted in Fig. 3, show a Gaussian distribution in the weak monitoring limit, consistent with the central limit theorem for small, uncorrelated measurement perturbations. As the monitoring strength increases, deviations from Gaussianity are gradually observed. The full distribution is well described by $\sim\exp(-x^{2}/(a|x|+b))$ with $a,b$ fitting parameters, which suggests that stronger monitoring induces large fluctuations leading to broader exponential (not Gaussian) tails. The strong monitoring limit is characterized by a peak around zero, consistent (fitting not shown) with a soft power-law singularity at this point which indicates that the entanglement dynamics starts to be dominated by the quantum Zeno effect. The tails are still well described by the same expression that interpolates between a Gaussian and an exponential. Fittings with a stretched exponential (not shown) function also provide a good description of the numerical data so further research would be needed to fully describe the deviations from Gaussianity in this region.
Figure 3: The probability distribution of $\delta S_{qsd}$ for the QSD protocol Eq. (8), employing Eq. (7) after the saturation time of the average EE. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 3 to 3. The distribution agrees well with a Gaussian distribution in the weak monitoring limit. For intermediate monitoring strength, the distribution is well described by a simple function interpolating from Gaussian, for small $\delta S_{qsd}$, to exponential, for large $\delta S_{qsd}$. In the strong monitoring limit, the exponential tails are still observed. However, for small $\delta S_{qsd}$ , the distribution is strongly peaked around zero which indicates the dominance of quantum Zeno effect and provides an upper bound for the area law in the EE. We do not observe any size dependence in our results, see the text for an explanation of this feature.
@@ IV The quantum jump measurement protocol
We now proceed to investigate the monitored dynamics using the quantum jump protocol where the jump operator ${\hat{L}}_{i}=\sqrt{\gamma}\hat{n}_{i}$, representing the observable being measured, is proportional to the local occupation number $\hat{n}_{i}$, see Appendix A.2 for definitions and additional technical details. In order to study the effects of quantum jumps to the entanglement entropy, we define
@@
Results for the probability distribution of $\Delta S_{qj}$, presented in Fig. 4, are quite different from those of the QSD protocol, because for the latter we are performing measurements on all sites at each time step. Even for weak monitoring rates, the distribution has a strong peak at zero that becomes sharper as the monitoring strength, or the system size $L$ increases. This means that many measurements have no effect on the EE and that those zero measurements become more important as the thermodynamic limit $L\to\infty$ is approached. These results are consistent with the analytical prediction [37] for the PM protocol that for $L\to\infty$, the EE is in the area law phase for any finite measuring rate.

### v4

# arXiv:2504.04071v4[quant-ph] 02 Mar 2026

@@ Abstract
For one-dimensional non-interacting complex fermions, we compute numerically the probability distribution of the change in the entanglement entropy (EE) after saturation, resulting from a single measurement of the occupation number by using different measurements protocols. In the thermodynamic limit, the system is in the area-law phase for any monitoring strength, however we can also study the distribution in the critical phase characterized by a logarithmic scaling of the EE with system size by considering a sufficiently weak monitoring strength. For a quantum state diffusion protocol, where the change in EE is defined between two consecutive time steps, the distribution, Gaussian for weak monitoring, gradually develops symmetric exponential tails. For strong monitoring, the core turns from Gaussian to strongly peaked at zero suggesting the dominance of quantum Zeno effect. For the quantum jump and the projective measurement protocols, we observe clear deviations from Gaussianity characterized by broader and asymmetric tails, exponential for positive values of the change, and a peak at zero, likely a precursor of Zeno effect, that increases with the system size and the monitoring strength. Intriguingly, the distribution is spatially inhomogeneous. For sites around the boundary separating the subsystems defining the EE, the distribution is close to Gaussian with a broad support while for the rest of sites has asymmetric exponential tails and a much narrower support. As the monitoring strength increases, the full distribution is controlled by the boundary sites.
@@ I Introduction
The quantum measurement problem has gained new impetus in recent times not only because the availability of the mentioned experimental measurement set-ups with a high level of control but also because for quantum computation, and other quantum technologies, it is fundamental to extract information from quantum states with a minimal level of disturbance on the system so that quantum coherence is preserved. This is a challenging task as repeated measurements can trigger [23, 24, 25, 26, 27, 21, 28, 29, 30, 31, 32, 33, 34] a measurement induced phase transitions (MIPT) in the entanglement entropy, and other quantum information observables non-linear in the density matrix, separating phases with different scaling with respect to the system size. Therefore, a high monitoring rate has the potential of destroying quantum features that may be necessary for practical applications. Although challenging because of the post-selection problem [35, 36, 37], the existence of MIPT has already [38, 39] been confirmed experimentally.
@@
The precise characterization of the MIPT in interacting many-body quantum systems are still being debated. However, for one dimensional free Dirac fermions, namely, non-relativistic complex fermions, with no disorder and subjected to either continuous [19] or projective [40] measurements there is now growing consensus, despite some initial contradicting results [41, 20], that the system stays in the area-law phase for any monitoring strength. For projective measurements [40], see also [42, 43], this conclusion was reached analytically by using the replica trick and mapping the problem onto a non-linear sigma model [44, 45, 46, 47]. However, a MIPT has been reported for Majorana fermions in one dimension [48], for Dirac fermions with long-range hopping [49, 50, 51], in higher dimensions [52] or including interactions [53, 54, 51, 55, 56, 57, 58]. Conditions for the existence of MIPT [59] has also been studied in the context of the Sachdev-Ye-Kitaev model [60, 61, 62, 63, 64] and also [65, 66, 67] its gravity dual which illustrates the broad appeal of these problems in rather different communities.
@@
All the above findings are based on the study of averaged (over quantum trajectories) quantities. However, it is unclear whether the average is a good indicator of the observable because of the importance of rare events or other features leading to a non-Gaussian distribution with sufficiently broad tails. However, the study of the full distribution of probability of observables like the entanglement entropy [68] or the local particle density and current [69] has just started. Even in the simplest case of a non-interacting fermionic system in one dimension, studied in Refs. [68, 69], it is not yet known the precise form of these distributions for the different measurement protocols, its system size dependence or its role in the characterization of the transitions.
@@ II Model and Method
$\hat{H}=J\sum_{i=1}^{L}\left[\hat{c}_{i}^{\dagger}\hat{c}_{i+1}+\hat{c}_{i+1}^{\dagger}\hat{c}_{i}\right],$ (1)
where $\hat{c}_{i}$ and $\hat{c}_{i}^{\dagger}$, $i=1,2,\cdots L$ are the annihilation and creation operators respectively for fermions at site $i$. $J$ is the hopping strength which for convenience is set to $J=1$. We impose periodic boundary conditions $\hat{c}_{i}=\hat{c}_{i+L},\hat{c}_{i}^{\dagger}=\hat{c}_{i+L}^{\dagger}$ with $L$ the lattice size which we set to be even. The Hamiltonian commutes with the particle number operator $\hat{N}=\sum_{i=1}^{L}\hat{n}_{i}$ where $\hat{n}_{i}\equiv\hat{c}^{\dagger}_{i}\hat{c}_{i}$ is the occupation number operator at site $i$ which has eigenvalues $0,1$. The eigenvalues of $\hat{N}$ are the total number of fermions $N=1,2,\ldots$. We consider the system at half-filling $N=L/2$. The initial state, termed Néel state, in an occupation number basis is $|\psi(t=0)\rangle=|101010\cdots\rangle$.
@@
(2). Quantum jump (QJ) protocol [10, 11, 12, 5, 13], a continuous monitoring protocol usually referred to as photo-detection in quantum optics where the measurement apparatus is always active and the detection, usually termed a quantum jump, occurs with a certain probability that depends on the monitoring strength $\gamma$. In our case, what is detected is an occupied site, $n_{i}=1$. We model the time evolution of the state by the quantum trajectory method [7, 8, 6, 9] characterized by the jump operator $\hat{L}_{i}=\sqrt{\gamma}\hat{n}_{i}$. Further details are in Appendix A.2.
@@
$|\psi(t)\rangle=\prod_{k=1}^{N}\left[\sum_{j=1}^{L}U_{jk}(t)\hat{c}_{j}^{\dagger}\right]|\text{vac}\rangle.$ (2)
@@
$D_{ij}\equiv\langle\psi(t)|\hat{c}_{i}^{\dagger}\hat{c}_{j}|\psi(t)\rangle$ (4)
@@
$|\psi(t)\rangle=\sum_{\{j\}}U_{j_{1},1}(t)U_{j_{2},2}(t)\cdots U_{j_{N},N}(t)\hat{c}^{\dagger}_{j_{1}}\hat{c}^{\dagger}_{j_{2}}\cdots\hat{c}^{\dagger}_{j_{N}}|\text{vac}\rangle=\sum_{\{j\}}\left[\prod_{i=1}^{N}U_{j_{i},i}(t)\hat{c}^{\dagger}_{j_{i}}\right]|\text{vac}\rangle$ (5)
then we substitute Eq. (5) into Eq. (4),
$D_{mn}=\sum_{\{i\},\{j\}}\biggl(\prod_{k=1}^{N}U_{i_{k},k}^{*}\biggr)\,\biggl(\prod_{l=1}^{N}U_{j_{l},l}\biggr)\,\langle\text{vac}|\hat{c}_{i_{N}}\cdots\hat{c}_{i_{1}}\hat{c}_{m}^{\dagger}\hat{c}_{n}\hat{c}_{j_{1}}^{\dagger}\cdots\hat{c}_{j_{N}}^{\dagger}|\text{vac}\rangle=\sum_{k=1}^{N}U^{*}_{mk}U_{nk}$ (6)
@@
Our primary focus is the entanglement entropy, $S=-\operatorname{Tr}(\hat{\rho}_{A}\ln\hat{\rho}_{A})$, where $\hat{\rho}_{A}$ is the reduced density matrix for the subsystem $A=\{1,2,\dots,\ell\}$ with $\ell\in\{1,2,\dots,L\}$. The correlation matrix $D^{A}$ in sub-system $A$ is given by the $\ell\times\ell$ upper-left block of $D$, with elements defined as $D^{A}_{ij}=\operatorname{Tr}(\hat{\rho}_{A}\,\hat{c}_{i}^{\dagger}\hat{c}_{j})$. Since $\hat{\rho}_{A}$ is Gaussian, Wick’s theorem allows us to express it in terms of the eigenvalues of $D^{A}$ [70, 71, 72, 73],
$S(\ell,t)=-\sum_{i=1}^{\ell}\left(\lambda_{i}(t)\ln(\lambda_{i}(t))+(1-\lambda_{i}(t))\ln(1-\lambda_{i}(t))\right)$ (7)
@@
Strictly speaking, rather than with the distribution of the full EE at those long times, our main interest is the distribution of the EE difference between the state right before and right after a measurement – for homodyne detection this corresponds to one time step. However, for the sake of completeness, we depict in Fig. 2 the probability distribution of the full EE in this long time region. As is observed, in the weak monitoring limit, the distribution is close to Gaussian for the three protocols. Deviations from Gaussianity, characterized by broad tails and a bimodal distribution are qualitatively similar in all protocols. Such features become evident only when the monitoring rate is sufficiently strong, which we believe corresponds to a region where the quantum Zeno effect [74, 75, 76, 77, 78] is dominant. We comment that similar results for the full EE distribution were obtained earlier for continuously monitored Ising [31] and free-fermion chains [20]. These results suggest that the distribution function of the full EE seems to be capable of capturing a transition in the EE (if present). As we mentioned in the introduction, this model is in the area-law phase for $L\to\infty$, and any $\gamma>0$, but we still believe that in the range of $L$ that we explore, and for small $\gamma$, our results describe the distribution of a system in a regime where the EE scales with the logarithmic of the system size. In any case, it is promising that within the range of sizes we can explore, the full distribution captures sharp differences between the weak and strong monitoring rates.
@@ III The quantum state diffusion measurement protocol
We initiate our analysis with the study of the distribution function for the change of EE for the case of continuous measurements characterized by the QSD protocol, see Appendix A.1 for details. We shall show that for this distribution, the distribution is Gaussian for weak monitoring. As monitoring strength increases, the tails becomes broader and the center develops a peak that suggests the growing importance of the Zeno effect. This protocol, related to homodyne measurements, is characterized by weak measurements of the particle number on each site of the lattice at each time step $dt$. Therefore, by change in entanglement after one measurement, we really mean the change in EE after one time step because, by definition, at a given time, measurements are carried out on all sites at once so the change after a single measurement is not possible. It is natural to study the speed of the entanglement change at each step $dt$ of the dynamical evolution defined as,
@@
where $t$ is sufficiently large so that the EE has already reached its saturation value and $S_{A}$ is computed using Eq. (7).
III.1 Main Results
The probability distribution of $\delta S_{qsd}$, depicted in Fig. 3, show a Gaussian distribution in the weak monitoring limit, consistent with the central limit theorem for small, uncorrelated measurement perturbations. As the monitoring strength increases, deviations from Gaussianity are gradually observed. The full distribution is well described by $\sim\exp(-x^{2}/(a|x|+b))$ with $a,b$ fitting parameters, which suggests that stronger monitoring induces large fluctuations leading to broader exponential (not Gaussian) tails. The strong monitoring limit is characterized by a peak around zero, consistent (fitting not shown) with a soft power-law singularity at this point which indicates that the entanglement dynamics starts to be dominated by the quantum Zeno effect. The tails are still well described by the same expression that interpolates between a Gaussian and an exponential. Fittings with a stretched exponential (not shown) function also provide a good description of the numerical data so further research would be needed to fully describe the deviations from Gaussianity in this region.
Figure 3: The probability distribution of $\delta S_{qsd}$ for the QSD protocol in Eq. (8), employing Eq. (7) after the saturation time of the average EE. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 3 to 3. The distribution agrees well with a Gaussian distribution in the weak monitoring limit. For intermediate monitoring strength, the distribution is well described by a simple function interpolating from Gaussian, for small $\delta S_{qsd}$, to exponential, for large $\delta S_{qsd}$. In the strong monitoring limit, the exponential tails are still observed. However, for small $\delta S_{qsd}$, the distribution is strongly peaked around zero which indicates the dominance of quantum Zeno effect. We do not observe any size dependence in our results, see the text for an explanation of this feature.
@@ IV The quantum jump measurement protocol
We now study the monitored dynamics employing the quantum jump protocol where the jump operator ${\hat{L}}_{i}=\sqrt{\gamma}\hat{n}_{i}$, representing the observable being measured, is proportional to the local occupation number $\hat{n}_{i}$, see Appendix A.2 for definitions and additional technical details. We shall see that the main differences with respect to the QSD protocol are that the distribution has a peak at zero even for weak monitoring, and that its tails for positive and negative changes in entropy are qualitatively different. Moreover, the change in entropy is quite sensitive to the site position. In order to study the effects of quantum jumps to the entanglement entropy, we define
@@
IV.1 Distribution of $\Delta S_{qj}$ after each measurement
Results for the probability distribution of $\Delta S_{qj}$, presented in Fig. 4, are quite different from those of the QSD protocol, because for the latter we are performing measurements on all sites at each time step. Even for weak monitoring rates, the distribution has a strong peak at zero that becomes sharper as the monitoring strength, or the system size $L$ increases. This means that many measurements have no effect on the EE and that those measurements become more important as the thermodynamic limit $L\to\infty$ is approached. These results are consistent with the analytical prediction [40] for the PM protocol that for $L\to\infty$, the EE is in the area law phase for any finite measuring rate.